opdaxy:

4. What is 25% of 80? * @dk4l @xiogonz

opdaxy:

ok

DK4L:

are there options?

jhonyy9:

80 is 100% x is 25% cross multiplie and calcule x x = 80*25/100 = ?

opdaxy:

5. What is 20% of 75? *

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @opdaxy 5. What is 20% of 75? * \(\color{#0cbb34}{\text{End of Quote}}\) this was before solved

XioGonz:

Here \[\frac{ 25*80 }{ 100 }=20\] Multiply 80 by the percent then divide by 100 Same applies to other Q's like this

opdaxy:

Three students wrote equations on their dry erase boards. Which of these students wrote an equation that is true? * A. Student A and Student C B. Student B and Student C C. Student B D. Student C